OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = -3(x + 8)2 + 5? x = -3 x = 5 x = 8 x = -8

OpenStudy (driftracer305):

can u graph it..??

OpenStudy (driftracer305):

@kellyconnelly ?

OpenStudy (anonymous):

is y = -3(x + 8)2 + 5? or y = -3(x + 8)^2 + 5?

OpenStudy (driftracer305):

it is squared........ she wont square it..... lol

OpenStudy (anonymous):

second one and ill square it >:(

OpenStudy (anonymous):

max value = 5 when x+8=0 so x=-8 is the x of the max pt, for quadratic, graph symmetry at max pt, so x=-8 is the answer

OpenStudy (driftracer305):

yea so what x value do u think will cut it exactly in half..?

OpenStudy (anonymous):

how u make graphs ? i mean software ?

OpenStudy (anonymous):

http://www.wolframalpha.com/ @ramit.dour

OpenStudy (anonymous):

actually this equation ald done the square completing, you can directly obtain answer in the equation without drawing graph

OpenStudy (driftracer305):

lol @neoh147 ....... yea u gave her the answer..... :D